OpenStudy (anonymous):

express in terms of i the sum of radical -25 + 2 (radical -36) I got 11i + 2 am I right???

OpenStudy (anonymous):

\[-25+2\sqrt{-36}=-25+2(6)\sqrt{-1}=-25+12i\]

OpenStudy (anonymous):

no it's radical negative 25

OpenStudy (anonymous):

so\[\sqrt{-25}+2\sqrt{-36}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{-25}+2\sqrt{-36}=5\sqrt{-1}+2(6)\sqrt{-1}=5i+12i=17i\]

OpenStudy (anonymous):

\[\sqrt{36}\] is 12i????

OpenStudy (anonymous):

\[\sqrt{-36}=\sqrt{-1*(6^2)}=6\sqrt{-1}=6i\] However, the equation that we worked had a coefficient of two so \[2\sqrt{-36}=12i\]

OpenStudy (anonymous):

ohhh okay I see, Thanks!!!